OpenStudy (anonymous):

how do I change the top to get the 20 by the equal sign? W +T=A+20 A+W+T=90

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

not exactly sure what you're asking

jimthompson5910 (jim_thompson5910):

do you want to move that A on the right side over to the left side?

OpenStudy (anonymous):

likef or example a+w+t=90 i would like to get the first equation similar like for example a+w+t=20

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

we have A+20 on the right side or 20+A

jimthompson5910 (jim_thompson5910):

to undo that +A, we have to subtract A from both sides

jimthompson5910 (jim_thompson5910):

that will give us W +T=A+20 W +T-A=A+20-A W +T-A=20 and you can rearrange terms to get -A+W+T = 20

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw